OpenStudy (anonymous):

Use this quation to find the average height of a 13 year old girl. y describes the height of a girl and x is her age y - 44 = 5(x-2.4) Select one: a. 63.2 in b. 56.3 in c. 69.4 in d. 97 in

OpenStudy (anonymous):

y-44=5(x-2.4) y-44=5(13-2.4) y-44=5(10.6) y-44=53 y=53+44 y= 97 in.